OpenStudy (anonymous):

A cylindrical (right circular cylinder) package to be sent by a postal service can have a maximum combined length and girth (perimeter of a cross section) of 108 inches. Find the dimensions of the package of maximum volume that can be sent. (assume the cross section is circular)

OpenStudy (anonymous):

I got V = 2916 in^3

OpenStudy (cwrw238):

L + 2pi R = 108 V = pi R^2 L substituting in V for L V = piR^2 (108 - 2Pi R) find dV / dR and equate to 0 to find R for maxm V is that the way you did it?

OpenStudy (anonymous):

Yes, precisely! I'm just not sure if my answer is right.

OpenStudy (cwrw238):

dV/dR = 216piR - 6pi^2 R^2 = 0 6piR (piR - 36) = 0 R = 36 / pi = 11.459 so L = 108 - 2 pi * 11.459 = 36.001 V = pi * 11.459^2 * 36.001 = 14851 if my calculations are correct!!

OpenStudy (anonymous):

Ah, okay. I think I messed up on the derivative part.

OpenStudy (cwrw238):

i've checked it out and cant find anything wrong

OpenStudy (anonymous):

So pi just stays as is when you take a derivative?

OpenStudy (cwrw238):

yea pi is a constant

OpenStudy (anonymous):

Blubber nuggets! Okay. Thanks!

OpenStudy (cwrw238):

lol - yw